### v1

# Minimisation of Laplacian eigenvalue with indefinite weight under inhomogeneous Robin boundary condition

Abstract
This paper explores a certain Laplacian eigenvalue optimisation problem with indefinite weight under inhomogeneous Robin boundary condition. The minimum principal eigenvalue is fully determined in one dimension by formulating the problem as a shape optimisation one. The result is verified numerically using a shooting method.
pacs:
MSC Classification]35P15, 35J25
1 Introduction
In this paper, we consider the following eigenvalue problem
$\left\{\begin{aligned} \Delta\phi+\lambda m\phi&=0&\ \text{in}&\ \Omega\\ \partial_{\mathbf{n}}\phi+\beta\phi&=0&\ \text{on}&\ \partial\Omega\end{aligned}\right$. (1)
where $\Omega$ is a bounded domain in $\mathbb{R}^{n}$ with Lipschitz boundary, $\mathbf{n}$ is the unit outer normal on $\partial\Omega$, and the weight function $m$ is a bounded measurable function that changes sign in $\Omega$, i.e. $\Omega_{m}^{+}:=\{x\in\Omega;\ m(x)>0\}$ satisfies $|\Omega_{m}^{+}|\in(0,|\Omega|)$, and that
$1\leq m(x)\leq\kappa\ \text{a.e.}\ x\in\Omega$
for a constant $\kappa>0$. The inhomogeneous Robin boundary condition applies, with $\beta:\partial\Omega\to\mathbb{R}$ a piecewise continuous function.
According to [1], [2] and [3], $\lambda$ is called a principal eigenvalue of (1), if the associated eigenfunction $\phi$ is positive, where $\phi\in H^{1}(\Omega)$.
The investigation of Laplacian eigenvalues with an indefinite weight under Robin boundary conditions originates from studies of diffusive logistics models in a bounded domain $\Omega$ with suitable boundary. In [4], it is shown that the density $u(x,t)$ of the species at location $x$ and time $t$ satisfies the following equation:
$\begin{cases}u_{t}=\Delta u+\omega u[m(x)-u]\ \text{in}\ \Omega\times\mathbb{R}^{+},\\ \partial_{\mathbf{n}}u+\beta_{0}u=0\ \text{on}\ \partial\Omega\times\mathbb{R}^{+},\\ u(x,0)\geq 0,\ u(x,0)\not\equiv 0\ \text{in}\ \bar{\Omega},\end{cases}$ (2)
where $\omega>0$ is a certain parameter. The Robin boundary parameter $\beta_{0}$ represents the inhospitableness of $\partial\Omega$. If $\beta_{0}=0$, the boundary is a barrier; if $\beta_{0}=+\infty$, the boundary is hostile. $m$ represents the intrinsic growth rate of the species, changing sign in $\Omega$, so that $\int_{\Omega}m$ represents the total resources in the spatially heterogeneous environment. In [5], the relation between the minimal positive principal eigenvalue $\lambda=\lambda(m)$ of (1) and the parameter $\omega$ in (2) is discovered: if $\omega\leq\lambda(m)$, then $u(x,t)\to 0$ uniformly in $\bar{\Omega}$ as $t\to\infty$ for proper initial data, i.e. the species extinguish; if $\omega>\lambda(m)$, then $u(x,t)\to u^{*}(x)$ uniformly in $\bar{\Omega}$ as $t\to\infty$, where $u^{*}$ is the unique positive steady solution of (2) in $W^{2,q}(\Omega)$ for every $q>1$, i.e. the species survives. Consequently, it is natural to consider the problem of minimising such an eigenvalue.
In this paper we consider, more generally, the inhomogeneous Robin boundary condition, and we focus on the case $\beta\geq 0$ on $\partial\Omega$.
We first prove the existence and uniqueness of a positive principal eigenvalue following the method by [6].
Lemma 1 ([7], Theorem 11.4 and 11.10).
For $\phi\in H^{1}(\Omega)$, let
$Q_{\phi}(\lambda):={\int_{\Omega}|\nabla\phi|^{2}+\int_{\partial\Omega}\beta\phi^{2}-\lambda\int_{\Omega}m\phi^{2}},$
and
$\mu(\lambda):=\inf_{\phi\in H^{1}(\Omega)}\frac{Q_{\phi}(\lambda)}{\int_{\Omega}\phi^{2}}.$
Then $\lambda$ is a principal eigenvalue of (1), if and only if $\mu(\lambda)=0$.
Lemma 2 ([6], Lemma 2).
If $\phi_{0}\in H^{1}(\Omega)$ is a minimiser of $Q_{\phi}(\lambda)$, i.e. $\mu(\lambda)=\frac{Q_{\phi_{0}}(\lambda)}{\int_{\Omega}\phi_{0}^{2}}$, then
$\mu^{\prime}(\lambda)=-\frac{\int_{\Omega}m\phi_{0}^{2}}{\int_{\Omega}\phi_{0}^{2}}.$
For any $\phi\in H^{1}(\Omega)$, the map $\lambda\mapsto Q_{\phi}(\lambda)$ is affine, hence concave. Since
$\mu(\lambda)=\inf\left\{Q_{\phi}(\lambda);\ \phi\in H^{1}(\Omega),\int_{\Omega}\phi^{2}=1\right\},$
the map $\mu:\lambda\mapsto\mu(\lambda)$ is also concave as the infimum of a collection of concave functions. In other words, $\mu^{\prime\prime}(\lambda)\leq 0$. Meanwhile, it is obvious that there exist $\phi_{1},\phi_{2}\in H^{1}(\Omega)$ satisfying $\int_{\Omega}m\phi_{1}^{2}>0$ and $\int_{\Omega}m\phi_{2}^{2}<0$, hence $\lim_{\lambda\to+\infty}\mu(\lambda)=\lim_{\lambda\to-\infty}\mu(\lambda)=$-$\infty$. Thus, $\mu=\mu(\lambda)$ increases until it attains its maximum, and decreases thereafter.
When $\beta>0$ on $\partial\Omega_{\beta}^{+}\subseteq\partial\Omega$ with $|\Omega_{\beta}^{+}|>0$, using the variational characterisation of the first Robin eigenvalue $\lambda_{1}(\beta)>0$, we have
$Q_{\phi}(0)\geq\lambda_{1}(\beta)\int_{\Omega}\phi^{2}>0$
whenever $\int_{\Omega}\phi^{2}>0$, hence $\mu(0)>0$, so $\mu$ has exactly two zeros, one positive and the other negative. In other words, (1) has a unique positive principal eigenvalue and a unique negative principal eigenvalue.
When $\beta\equiv 0$ a.e. on $\partial\Omega$, the situation is similar to Neumann boundary condition, and by taking the constant $\phi_{c}\equiv c$, we see
$Q_{\phi_{c}}(0)=0+c^{2}\int_{\partial\Omega}\beta=0,$
and thus $\mu(0)=0$. Since the associated eigenfunction is clearly a constant, by Lemma 2, we get:
  • if $\int_{\Omega}m<0$, then $\mu^{\prime}(0)>0$, and $\mu$ has a unique positive zero corresponding to the unique positive principal eigenvalue of (1);
  • if $\int_{\Omega}m>0$, then $\mu^{\prime}(0)<0$, and $\mu$ has a unique negative zero corresponding to the unique negative principal eigenvalue of (1);
  • if $\int_{\Omega}m>0$, then $\mu^{\prime}(0)=0$, and $\mu$ has no other zeros, hence (1) has a unique principal eigenvalue 0.
It is also well-known (cf. [8], [9], [1] and [3]) that the eigenvalue can be expressed with a Rayleigh quotient. To sum up, we have:
Theorem 3.
If $\beta\geq 0$ on $\partial\Omega$ and $\beta>0$ on a subset of positive measure, the problem (1) admits a unique positive principal eigenvalue $\lambda$, depending on $m$. Moreover,
$\lambda(m)=\inf_{\phi\in\mathcal{S}(m)}\frac{\int_{\Omega}|\nabla\phi|^{2}+\int_{\partial\Omega}\beta\phi^{2}}{\int_{\Omega}m\phi^{2}},$
where $\mathcal{S}(m):=\left\{\phi\in H^{1}(\Omega);\int_{\Omega}m\phi^{2}>0\right\}$.
Principal eigenvalues also admit monotonicity [^1] We remark that an attempt to derive strict monotonicity in [10], Theorem 2.4, fails, for a positive eigenfunction could vanish on the boundary. with respect to Robin parameter $\beta$.
Theorem 4.
Suppose $\lambda_{1},\lambda_{2}$ are principal eigenvalues of problem (1) with Robin parameters $\beta_{1}$, $\beta_{2}$, respectively. And denote by $\phi_{1}$ and $\phi_{2}$ the positive eigenfunctions associated with $\lambda_{1}$ and $\lambda_{2}$, respectively. Assuming $\int_{\Omega}m\phi_{1}^{2}>0$, if $\beta_{1}\geq\beta_{2}$, then $\lambda_{1}\geq\lambda_{2}$. If, in addition, we suppose that $\phi_{1}$ does not vanish on $\partial\Omega$ and $\beta_{1}>\beta_{2}$, then $\lambda_{1}>\lambda_{2}$.
Proof.
By Lemma 1,
$\int_{\Omega}|\nabla\phi_{1}|^{2}+\int_{\partial\Omega}\beta_{1}\phi_{1}^{2}-\lambda_{1}\int_{\Omega}m\phi_{1}^{2}$
$=0$
$=\inf_{\phi\in H^{1}(\Omega)}\left(\int_{\Omega}|\nabla\phi|^{2}+\int_{\partial\Omega}\beta_{2}\phi^{2}-\lambda_{2}\int_{\Omega}m\phi^{2}\right)$
$\leq\int_{\Omega}|\nabla\phi_{1}|^{2}+\int_{\partial\Omega}\beta_{2}\phi_{1}^{2}-\lambda_{2}\int_{\Omega}m\phi_{1}^{2},$
so
$\lambda_{1}-\lambda_{2}\geq\frac{\int_{\partial\Omega}(\beta_{1}-\beta_{2})\phi_{1}^{2}}{\int_{\Omega}m\phi_{1}^{2}}\geq 0.$
We are interested in the minimisation of the positive principal eigenvalue $\lambda(m)$:
$\inf_{m\in\mathcal{M}_{m_{0},\kappa}}$ $\lambda(m),$ (3)
$\mathcal{M}_{m_{0},\kappa}$ $=\left\{m\in L^{\infty}(\Omega);\ -1\leq m\leq\kappa,\left|\Omega_{m}^{+}\right|>0,\int_{\Omega}m\leq-m_{0}|\Omega|\right\}$
Here $m_{0}$ is a real constant such that $m_{0}\in(-\kappa,1)$ if $\beta>0$ and $m_{0}\in(0,1)$ if $\beta=0$.
The following result is proven in [1] for Neumann boundary condition ($\beta\equiv 0$), but the extension to general Robin case is straightforward:
Theorem 5.
There exists some weight function $m^{*}\in\mathcal{M}_{m_{0},\kappa}$ such that the infimum of (3) is attained at $\lambda(m^{*})$. Moreover, there exists a measurable subset $E^{*}$ of $\Omega$ such that
$m^{*}=\kappa\chi_{E^{*}}-\chi_{\Omega\setminus E^{*}}\text{ a.e. }x\in\Omega,$ (4)
where $\chi_{U}$ denotes the characteristic function of a set $U$. In addition, the volume constraint is active at $m^{*}$:
$\int_{\Omega}m^{*}=-m_{0}|\Omega|.$
In other words, the optimal weight function $m^{*}$ satisfies a bang-bang distribution in $\Omega$ with respect to some optimal set $E^{*}$. The problem can therefore be formulated as a shape optimisation one:
$\inf_{E\in\mathcal{E}_{c,\kappa}}\lambda(E)$ $=\lambda(\kappa\chi_{E}-\chi_{\Omega\setminus E}),$ (5)
$\mathcal{E}_{c,\kappa}$ $=\{E\subseteq\Omega;E\ \text{is measurable},|E|\in(0,c|\Omega|]\}$
where $c:=\frac{1-m_{0}}{1+\kappa}$ and $c\in(0,1)$ if $\beta>0$ while $c\in\left(0,\frac{1}{1+\kappa}\right)$ if $\beta=0$. The optimal $E$ is expressed as a function of $\beta$, $\kappa$ and $c$.
In [2] and [3], this minimisation problem, formulated as a shape optimisation one, is thoroughly investigated under the usual homogeneous Robin boundary condition, with the 1-dimensional case fully solved. Unfortunately, for higher dimensional cases, few results are known. In [2], they studied this problem in a cylindrical domain with Robin boundary condition on the top and bottom and Neumann boundary condition on the lateral boundary. In [3], it is shown that when $\partial\Omega$ is connected, the optimal set is not always a ball. In [11], the 1-dimensional result is generalised onto spherical shells of any dimension.
2 The 1-dimensional case
In this section, we derive the complete solution of the problem (1) in one dimension.
Consider the problem (1) in $\Omega=(0,1)\subseteq\mathbb{R}$. The following result is shown in [3] for homogeneous Robin conditions, but its generalisation is straightforward.
Proposition 6.
If $\Omega=(0,1)$, then any optimal set $E^{*}$ for the problem (5) is an interval.
Now let the optimal set be the interval $E=(a,b)$ where $0<a<b<1$ and c=bac=b-a. We intend to find the optimal value of $a$ following the method by [2]. Let $\beta_{0}=\beta(0)\geq 0$ and $\beta_{1}=\beta(1)>0$. Rewrite the problem using $m=\kappa\chi_{E}-\chi_{\Omega\setminus E}$, and we have
$\left\{\begin{aligned} \phi^{\prime\prime}-\lambda\phi&=0&\ \text{in}&\ (0,a),\\ \phi^{\prime\prime}+\lambda\kappa\phi&=0&\ \text{in}&\ (a,b),\\ \phi^{\prime\prime}-\lambda\phi&=0&\ \text{in}&\ (b,1),\\ \phi^{\prime}(0)-\beta_{0}\phi(0)&=0,\\ \phi^{\prime}(1)+\beta_{1}\phi(1)&=0,\\ \phi(a)^{+}-\phi(a)^{-}&=0,\\ \phi(b)^{+}-\phi(b)^{-}&=0,\\ \phi^{\prime}(a)^{+}-\phi^{\prime}(a)^{-}&=0,\\ \phi^{\prime}(b)^{+}-\phi^{\prime}(b)^{-}&=0.\end{aligned}\right$. (6)
Since $\lambda>0$, $\phi>0$, we may assume
$\phi(x)=\left\{\begin{aligned} C_{1}\cosh\left(\sqrt{\lambda}(x-a)\right)+C_{2}\sinh\left(\sqrt{\lambda}(x-a)\right)\ \text{in}\ (0,a),\\ C_{3}\cos\left(\sqrt{\lambda\kappa}(x-a)\right)+C_{4}\sin\left(\sqrt{\lambda\kappa}(x-a)\right)\ \text{in}\ (a,b),\\ C_{5}\cosh\left(\sqrt{\lambda}(x-b)\right)+C_{6}\sinh\left(\sqrt{\lambda}(x-b)\right)\ \text{in}\ (b,1),\end{aligned}\right$. (7)
where $C_{i}=C_{i}(E)$ are constants for $1\leq i\leq 6$. Now we simplify the boundary and transmission conditions in (6). By organising the coefficients of $C_{i}$ ($1\leq i\leq 6$), we get
$\left\{\begin{aligned} &\left(-\sqrt{\lambda}\sinh(a\sqrt{\lambda})-\beta_{0}\cosh(a\sqrt{\lambda})\right)C_{1}+\left(\sqrt{\lambda}\cosh(a\sqrt{\lambda})+\beta_{0}\sinh(a\sqrt{\lambda})\right)C_{2}=0\\ &\left(\sqrt{\lambda}\sinh(\sqrt{\lambda}(1-b))+\beta_{1}\cosh(\sqrt{\lambda}(1-b))\right)C_{5}\\ &\ \ \ +\left(\sqrt{\lambda}\cosh(\sqrt{\lambda}(1-b))+\beta_{1}\sinh(\sqrt{\lambda}(1-b))\right)C_{6}=0\\ &C_{3}-C_{1}=0\\ &C_{5}-C_{3}\cos(\sqrt{\lambda\kappa}c)-C_{4}\sin(\sqrt{\lambda\kappa}c)=0\\ &\sqrt{\kappa}C_{4}-C_{2}=0\\ &C_{6}+C_{3}\sqrt{\kappa}\sin(\sqrt{\lambda\kappa}c)-C_{4}\sqrt{\kappa}\cos(\sqrt{\lambda\kappa}c)=0\end{aligned}\right$. (8)
Using the last 4 equations, we can express $C_{1},C_{2},C_{5},C_{6}$ in terms of $C_{3}$ and $C_{4}$, and then we obtain two linear equations of $C_{3}$ and $C_{4}$ only:
$\left\{\begin{aligned} &\left(-\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(a)-\beta_{0}{\mathfrak{c}}_{\lambda}(a)\right)C_{3}+\sqrt{\kappa}\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(a)+\beta_{0}{\mathfrak{s}}_{\lambda}(a)\right)C_{4}=0\\ &\left[\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)-\sqrt{\kappa}\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right]C_{3}\\ &+\left[\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)+\sqrt{\kappa}\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right]C_{4}=0\end{aligned}\right$. (9)
where ${\mathfrak{s}}_{\lambda}(a):=\sinh(\sqrt{\lambda}a)$, ${\mathfrak{c}}_{\lambda}(a):=\cosh(\sqrt{\lambda}a)$, ${\mathfrak{s}}_{\lambda}(b):=\sinh(\sqrt{\lambda}(1-b))$ and ${\mathfrak{c}}_{\lambda}(b):=\cosh(\sqrt{\lambda}(1-b))$. Considered as a $2\times 2$ matrix applied to the column vector $(C_{3},C_{4})^{t}$, they have a non-trivial solution if and only if the determinant of the matrix vanishes, i.e.
$\left(-\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(a)-\beta_{0}{\mathfrak{c}}_{\lambda}(a)\right)\left[\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)\right.\\ +\left.\sqrt{\kappa}\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right]\\ =\sqrt{\kappa}\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(a)+\beta_{0}\sinh(a\sqrt{\lambda})\right)\left[\cos(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{s}}_{\lambda}(b)+\beta_{1}{\mathfrak{c}}_{\lambda}(b)\right)\right.\\ -\left.\sqrt{\kappa}\sin(\sqrt{\lambda\kappa}c)\left(\sqrt{\lambda}{\mathfrak{c}}_{\lambda}(b)+\beta_{1}{\mathfrak{s}}_{\lambda}(b)\right)\right].$ (10)
First we order the term and get
$\cosh(\sqrt{\lambda}(1-b))\left[\cosh(\sqrt{\lambda}a)\left(-(\beta_{0}+\beta_{1})\sqrt{\lambda\kappa}\cos(\sqrt{\lambda\kappa}c)+(\lambda\kappa-\beta_{0}\beta_{1})\sin(\sqrt{\lambda\kappa}c)\right)\right.\\ \left.+\sinh(\sqrt{\lambda}a)\left(-\sqrt{\kappa}(\lambda+\beta_{0}\beta_{1})\cos(\sqrt{\lambda\kappa}c)+(\kappa\beta_{0}-\beta_{1})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)\right)\right]\\ +\sinh(\sqrt{\lambda}(1-b))\left[\cosh(\sqrt{\lambda}a)\left(-\sqrt{\kappa}(\lambda+\beta_{0}\beta_{1})\cos(\sqrt{\lambda\kappa}c)+(\kappa\beta_{1}-\beta_{0})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)\right)\right.\\ \left.+\sinh(\sqrt{\lambda}a)\left(-\sqrt{\lambda\kappa}(\beta_{0}+\beta_{1})\cos(\sqrt{\lambda\kappa}c)+(\beta_{0}\beta_{1}\kappa-\lambda)\sin(\sqrt{\lambda\kappa}c)\right)\right]=0.$ (11)
Let us first verify (11) in two extreme cases. When $\beta_{0}=\beta_{1}=0$, the problem is reduced to Neumann boundary condition, and we have obtained the characteristic equation [^2] We remark that the characteristic equations (18) and (20) obtained in [2] based on homogeneous Robin boundary condition admit the same typo on the left-hand side.
$\tanh(\lambda^{\frac{1}{2}}(1-b))=\frac{\kappa^{\frac{1}{2}}\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)-\tanh(\lambda^{\frac{1}{2}}a)}{1+\kappa^{-\frac{1}{2}}\tanh(\lambda^{\frac{1}{2}}a)\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)},$ (12)
which reduces, when $a=0$, to the equation (3.3) in [1]:
$\kappa^{\frac{1}{2}}\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)=\tanh(\lambda^{\frac{1}{2}}(1-c)).$ (13)
On the other hand, when $\beta_{0},\beta_{1}\to+\infty$, the problem is reduced to Dirichlet boundary condition, we similarly get the characteristic equation
$\tanh(\lambda^{\frac{1}{2}}(1-b))=\frac{\kappa^{-\frac{1}{2}}\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)+\tanh(\lambda^{\frac{1}{2}}a)}{\kappa^{\frac{1}{2}}\tanh(\lambda^{\frac{1}{2}}a)\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)-1},$ (14)
which reduces, when $a=0$, to the equation (21) in [2]:
$\tan(\lambda^{\frac{1}{2}}\kappa^{\frac{1}{2}}c)=-\kappa^{\frac{1}{2}}\tanh(\lambda^{\frac{1}{2}}(1-c)).$ (15)
Now we use the product-to-sum formulae for hyperbolic functions to simplify it. Using $a\sqrt{\lambda}+\sqrt{\lambda}(1-b)=\sqrt{\lambda}(1-c)$ we get
$0=(\kappa+1)(\lambda-\beta_{0}\beta_{1})\cosh(\sqrt{\lambda}(a+b-1))\sin(\sqrt{\lambda\kappa}c)\\ +(\kappa+1)(\beta_{0}-\beta_{1})\sqrt{\lambda}\sinh(\sqrt{\lambda}(a+b-1))\sin(\sqrt{\lambda\kappa}c)\\ +\cosh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\lambda+\beta_{0}\beta_{1})\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\lambda\kappa}(\beta_{0}+\beta_{1})\cos(\sqrt{\lambda\kappa}c)\right)\\ +\sinh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\beta_{0}+\beta_{1})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\kappa}(\beta_{0}\beta_{1}+\lambda)\cos(\sqrt{\lambda\kappa}c)\right).$ (16)
We express the right-hand side of (16) as a function $f(a,\beta_{0},\beta_{1},\kappa,\lambda)$, using $b=a+c$:
$f(a,\beta_{0},\beta_{1},\kappa,\lambda)=(\kappa+1)(\lambda-\beta_{0}\beta_{1})\cosh(\sqrt{\lambda}(2a+c-1))\sin(\sqrt{\lambda\kappa}c)\\ +(\kappa+1)(\beta_{0}-\beta_{1})\sqrt{\lambda}\sinh(\sqrt{\lambda}(2a+c-1))\sin(\sqrt{\lambda\kappa}c)\\ +\cosh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\lambda+\beta_{0}\beta_{1})\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\lambda\kappa}(\beta_{0}+\beta_{1})\cos(\sqrt{\lambda\kappa}c)\right)\\ +\sinh(\sqrt{\lambda}(1-c))\left((\kappa-1)(\beta_{0}+\beta_{1})\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)-2\sqrt{\kappa}(\beta_{0}\beta_{1}+\lambda)\cos(\sqrt{\lambda\kappa}c)\right).$ (17)
Taking partial derivative with respect to $a$ gives
$\partial_{a}f(a,\beta_{0},\beta_{1},\kappa,\lambda)=2(\kappa+1)\sqrt{\lambda}\sin(\sqrt{\lambda\kappa}c)\cosh(\sqrt{\lambda}(2a+c-1))\\ \cdot\left[(\lambda-\beta_{0}\beta_{1})\tanh\left(2\sqrt{\lambda}\left(a-\frac{1-c}{2}\right)\right)+\sqrt{\lambda}(\beta_{0}-\beta_{1})\right].$ (18)
Since we are concerned with the first eigenvalue only, we assume $0<\lambda<\frac{\pi^{2}}{4c^{2}\kappa}$, so $\sin(\sqrt{\lambda\kappa}c)>0$ and $\cos(\sqrt{\lambda\kappa}c)>0$. Therefore, the sign of $\partial_{a}f$ depends on that of
$g(a,\beta_{0},\beta_{1},\lambda):=(\lambda-\beta_{0}\beta_{1})\tanh\left(2\sqrt{\lambda}\left(a-\frac{1-c}{2}\right)\right)+\sqrt{\lambda}(\beta_{0}-\beta_{1}).$ (19)
Lemma 7.
For any $a\in[0,1-c]$, $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)\not\equiv 0$, where $\kappa>0$.
Proof.
Write $\partial_{\beta_{1}}f=A(a)\,\beta_{0}+B(a)$, and assume that for some $a\in[0,1-c]$ we have $A(a)=B(a)=0$, where $A,B$ are defined in (37). Then $A(a)=B(a)=0$ gives
$(\kappa+1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(2a+c-1)\big)$ = $=-\,2\sqrt{\kappa}\,\cos\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(1-c)\big)$
$\quad+(\kappa-1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(1-c)\big),$ (20)
$(\kappa+1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(2a+c-1)\big)$ $=(\kappa-1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(1-c)\big)$
$\quad-2\sqrt{\kappa}\,\cos\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(1-c)\big).$ (21)
Now square both (20) and (21) and subtract them. Using the identity $\cosh^{2}\!\big(\sqrt{\lambda}(2a+c-1)\big)-\sinh^{2}\!\big(\sqrt{\lambda}(2a+c-1)\big)=1$, we obtain
$(\kappa+1)^{2}\sin^{2}\!\big(c\sqrt{\kappa\lambda}\big)=P^{2}-Q^{2},$ (22)
where
$P$ = $=-\,2\sqrt{\kappa}\,\cos\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(1-c)\big)+(\kappa-1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(1-c)\big),$ (23)
$Q$ $=(\kappa-1)\,\sin\!\big(c\sqrt{\kappa\lambda}\big)\,\sinh\!\big(\sqrt{\lambda}(1-c)\big)-2\sqrt{\kappa}\,\cos\!\big(c\sqrt{\kappa\lambda}\big)\,\cosh\!\big(\sqrt{\lambda}(1-c)\big).$ (24)
Expanding $P^{2}-Q^{2}$,
$P^{2}-Q^{2}=\left[4\kappa\cos^{2}(c\sqrt{\kappa\lambda})\sinh^{2}(\sqrt{\lambda}(1-c))+(\kappa-1)^{2}\sin^{2}(c\sqrt{\kappa\lambda})\cosh^{2}(\sqrt{\lambda}(1-c))\right.\\ \left.-4\sqrt{\kappa}(\kappa-1)\cos(c\sqrt{\kappa\lambda})\sin(c\sqrt{\kappa\lambda})\cosh(\sqrt{\lambda}(1-c))\sinh(\sqrt{\lambda}(1-c))\right]\\ -\left[(\kappa-1)^{2}\sin^{2}(c\sqrt{\kappa\lambda})\sinh^{2}(\sqrt{\lambda}(1-c))+4\kappa\cos^{2}(c\sqrt{\kappa\lambda})\cosh^{2}(\sqrt{\lambda}(1-c))\right.\\ -\left.-4\sqrt{\kappa}(\kappa-1)\cos(c\sqrt{\kappa\lambda})\sin(c\sqrt{\kappa\lambda})\cosh(\sqrt{\lambda}(1-c))\sinh(\sqrt{\lambda}(1-c))\right]\\ =(\kappa-1)^{2}\sin^{2}(c\sqrt{\kappa\lambda})-4\kappa\cos^{2}(c\sqrt{\kappa\lambda})\\ =(\kappa+1)^{2}\sin^{2}(c\sqrt{\kappa\lambda})-4\kappa,$ (25)
and comparing (22) and (25) we get $\kappa=0$, a contradiction to the assumption that $\kappa>0$. ∎
Lemma 8.
Suppose $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, where
$c^{*}=\frac{1}{1+\frac{2\sqrt{\kappa}}{\pi}\ln\left(\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}\right)}.$ (26)
Then the function
$h(c,\kappa,\lambda)=\frac{(\kappa-1)\sin(c\sqrt{\kappa\lambda})\cosh(\sqrt{\lambda}(1-c))-2\sqrt{\kappa}\cos(c\sqrt{\kappa\lambda})\sinh(\sqrt{\lambda}(1-c))}{(\kappa+1)\sin(c\sqrt{\kappa\lambda})}$ (27)
satisfies $h<1$ for any $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$.
Proof.
Since $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$, $\sin(c\sqrt{\kappa\lambda})>0$ and $\cos(c\sqrt{\kappa\lambda})>0$, so we have
$h=\frac{\kappa-1}{\kappa+1}\cosh(\sqrt{\lambda}(1-c))-\frac{2\sqrt{\kappa}}{\kappa+1}\frac{\sinh(\sqrt{\lambda}(1-c))}{\tan(c\sqrt{\kappa\lambda})}.$ (28)
Let $t=c\sqrt{\kappa\lambda}\in\left(0,\frac{\pi}{2}\right)$ and $\theta=\sqrt{\lambda}(1-c)\in\left(0,\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)$, then
$h=h(\kappa,t,\theta)=\frac{\kappa-1}{\kappa+1}\cosh(\theta)-\frac{2\sqrt{\kappa}}{\kappa+1}\frac{\sinh(\theta)}{\tan(t)}<\frac{\kappa-1}{\kappa+1}\cosh(\theta)<\frac{\kappa-1}{\kappa+1}\cosh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right).$ (29)
Therefore, to ensure $h<1$, it suffices that
$\frac{\kappa-1}{\kappa+1}\cosh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)<1.$ (30)
Clearly, if $\kappa\leq 1$, (30) is satisfied automatically; if $\kappa>1$, (30) holds if and only if
$c>\frac{1}{1+\frac{2\sqrt{\kappa}}{\pi}\ln\left(\frac{\sqrt{\kappa}+1}{\sqrt{\kappa}-1}\right)}.$ (31)
We now prove:
Proposition 9.
Suppose $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, where $c^{*}$ is given in (26). Then there exists some $\beta_{0}^{*}=\beta_{0}^{*}(a)$ (given in (38)), such that $g(a,\beta_{0},\beta_{1},\lambda)$ and $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)$ have opposite signs if $\beta_{0}>\beta_{0}^{*}$, and they share the same sign if $\beta_{0}<\beta_{0}^{*}$.
Proof.
We consider $\lambda=\lambda(a,\beta_{0},\beta_{1},\kappa)$ as implicitly defined by $f(a,\beta_{0},\beta_{1},\kappa,\lambda)=0$. Differentiate by $a$, and we get
$\partial_{\lambda}f(a,\beta_{0},\beta_{1},\kappa,\lambda)\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)=-\partial_{a}f(a,\beta_{0},\beta_{1},\kappa,\lambda).$ (32)
Similarly, differentiate by $\beta_{1}$ and we get
$\partial_{\lambda}f(a,\beta_{0},\beta_{1},\kappa,\lambda)\partial_{\beta_{1}}\lambda(a,\beta_{0},\beta_{1},\kappa)=-\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda).$ (33)
It is shown in Theorem 4 that $\partial_{\beta_{0}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$ and $\partial_{\beta_{1}}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$. Now we have the implications:
$\partial_{\beta_{1}}f>0\Rightarrow\partial_{\lambda}f<0\Rightarrow\partial_{a}f\cdot\partial_{a}\lambda>0\Leftrightarrow g\cdot\partial_{a}\lambda>0,$ (34)
$\partial_{\beta_{1}}f<0\Rightarrow\partial_{\lambda}f>0\Rightarrow\partial_{a}f\cdot\partial_{a}\lambda<0\Leftrightarrow g\cdot\partial_{a}\lambda<0,$ (35)
$\partial_{\beta_{1}}f=0\Rightarrow\partial_{\lambda}f=0\Rightarrow\partial_{a}f=0\Leftrightarrow g=0.$ (36)
Notice that $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)=A(a)\beta_{0}+B(a)$ is a linear function with respect to $\beta_{0}$, where
$A(a)=-2\sqrt{\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)+(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)\\ -(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(2a+c-1)\big),\\[5.69054pt] B(a)=-2\sqrt{\lambda\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+(\kappa-1)\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)\\ -(\kappa+1)\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(2a+c-1)\big).$ (37)
We prove in Lemma 7 that $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)$ cannot vanish for any $a\in[0,1-c]$. Therefore, for each $a$, it has a unique zero
$\beta_{0}=\beta_{0}^{*}(a):=-\frac{B(a)}{A(a)}.$ (38)
Moreover, when $\beta_{0}>\beta_{0}^{*}(a)$ or $\beta_{0}<\beta_{0}^{*}(a)$, the sign of $\partial_{\beta_{1}}f(a,\beta_{0},\beta_{1},\kappa,\lambda)$ does not change, i.e. negative on one side and positive on the other. We now decide the monotonicity of $\partial_{\beta_{1}}f$ depending on $a$. The condition $A(a)>0$ is equivalent to
$\cosh(\sqrt{\lambda}(2a+c-1))<\frac{(\kappa-1)\sin(c\sqrt{\kappa\lambda})\cosh(\sqrt{\lambda}(1-c))-2\sqrt{\kappa}\cos(c\sqrt{\kappa\lambda})\sinh(\sqrt{\lambda}(1-c))}{(\kappa+1)\sin(c\sqrt{\kappa\lambda})}.$ (39)
Denote by $h=h(c,\kappa,\lambda)$ the expression on the right-hand side of (39). We show in Lemma 8 that $h<1$ for any $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$ under the assumed constraints on $c$, so $A(a)<0$ for any $a\in[0,1-c]$. In other words, $\partial_{\beta_{1}}f<0$ if $\beta_{0}>\beta_{0}^{*}$ and $\partial_{\beta_{1}}f>0$ if $\beta_{0}<\beta_{0}^{*}$. ∎
Lemma 10.
Denote by
$\tilde{\beta_{0}^{*}}(c,\kappa):=\frac{\sqrt{\kappa}\pi}{c}\frac{\sinh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)}{\kappa+1-(\kappa-1)\cosh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)},$ (40)
where $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, and $c^{*}$ is as in (26). Then $\beta_{0}^{*}(a)$ defined in (38) has an upper bound $\beta_{0}^{*}\leq\tilde{\beta_{0}^{*}}$, not depending on $a$ or $\lambda$.
Proof.
Notice that
$A(a)\geq 2\sqrt{\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)-(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)\\ +(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)\\ \geq-(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)>0,$ (41)
where the last inequality is guaranteed by the constraints on $c$, as shown in Lemma 8. Meanwhile,
$B(a)\leq B(0)=-2\sqrt{\lambda\kappa}\cos\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+2\kappa\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)\\ \leq 2\kappa\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)$ (42)
hence we have
$\beta_{0}^{*}$ $\leq\frac{2\kappa\sqrt{\lambda}\sin\big(c\sqrt{\kappa\lambda}\big)\sinh\big(\sqrt{\lambda}(1-c)\big)}{-(\kappa-1)\sin\big(c\sqrt{\kappa\lambda}\big)\cosh\big(\sqrt{\lambda}(1-c)\big)+(\kappa+1)\sin\big(c\sqrt{\kappa\lambda}\big)}$ (43)
$=\frac{2\kappa\sqrt{\lambda}\sinh\big(\sqrt{\lambda}(1-c)\big)}{(\kappa+1)-(\kappa-1)\cosh\big(\sqrt{\lambda}(1-c)\big)}.$ (44)
Notice that this upper bound is monotone increasing when $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$, hence
$\beta_{0}^{*}\leq\frac{\sqrt{\kappa}\pi}{c}\frac{\sinh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)}{\kappa+1-(\kappa-1)\cosh\left(\frac{\pi(1-c)}{2c\sqrt{\kappa}}\right)},$ (45)
where the right-hand side is a positive bound by the constraints on $c$ as shown in Lemma 8. ∎
Now we are ready to present the complete result on the 1-dimensional interval.
Theorem 11.
Consider problem (5) in $\Omega=(0,1)\subseteq\mathbb{R}^{1}$ with $\beta_{0}>\tilde{\beta_{0}^{*}}$, and $c\in(0,1)$ when $0<\kappa\leq 1$ and $c\in(c^{*},1)$ when $\kappa>1$, where $c^{*}$ is as in (26) and $\tilde{\beta_{0}^{*}}$ is as in (40). Then we have the following results. When $\beta_{0}\beta_{1}=\lambda$,
  • if $\beta_{0}>\beta_{1}$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)<0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for a=1ca=1-c;
  • if $\beta_{0}<\beta_{1}$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)>0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for $a=0$;
  • if $\beta_{0}=\beta_{1}=\lambda^{\frac{1}{2}}$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)=0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for any $a\in[0,1-c]$.
Otherwise, denote by
$a^{*}=\frac{1}{2\sqrt{\lambda}}\tanh^{-1}\left(\frac{\sqrt{\lambda}(\beta_{1}-\beta_{0})}{\lambda-\beta_{0}\beta_{1}}\right)+\frac{1-c}{2}.$ (46)
When $\beta_{0}\beta_{1}>\lambda$,
  • if $\beta_{0}-\beta_{1}<-\lambda^{-\frac{1}{2}}(\beta_{0}\beta_{1}-\lambda)\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\geq 0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for $a=0$;
  • if $|\beta_{0}-\beta_{1}|\leq\lambda^{-\frac{1}{2}}(\beta_{0}\beta_{1}-\lambda)\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\leq 0$ for $a\in\left(0,a^{*}\right)$, and $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\geq 0$ for $a\in\left(a^{*},1-c\right)$, so the minimum eigenvalue is attained at $a=a^{*}$;
  • if $\beta_{0}-\beta_{1}>\lambda^{-\frac{1}{2}}(\beta_{0}\beta_{1}-\lambda)\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\leq 0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for a=1ca=1-c.
When $\beta_{0}\beta_{1}<\lambda$,
  • if $\beta_{0}-\beta_{1}<-\lambda^{-\frac{1}{2}}(\lambda-\beta_{0}\beta_{1})\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\geq 0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for $a=0$;
  • if $|\beta_{0}-\beta_{1}|\leq\lambda^{-\frac{1}{2}}(\lambda-\beta_{0}\beta_{1})\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\geq 0$ for $a\in(0,a^{*})$, and $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\leq 0$ for $a\in(a^{*},1-c)$, so the minimum eigenvalue is attained at either $a=0$ or a=1ca=1-c;
  • if $\beta_{0}-\beta_{1}>\lambda^{-\frac{1}{2}}(\lambda-\beta_{0}\beta_{1})\tanh(\lambda^{\frac{1}{2}}(1-c))$, then $\partial_{a}\lambda(a,\beta_{0},\beta_{1},\kappa)\leq 0$ for $a\in(0,1-c)$, and the minimum eigenvalue is attained for a=1ca=1-c.
Proof.
The minimum of $\lambda$ can only be assumed at $a=0$, a=1ca=1-c or a possible internal minimal point, where $\partial_{a}\lambda=0$. By (32), a necessary condition satisfied by such an internal minimal point is $\partial_{a}f=0$. By (18) and (19), this happens if and only if $g=0$. First, suppose $\lambda=\beta_{0}\beta_{1}$. Then we see that if $\beta_{0}>\beta_{1}$, then $g>0$ for $a\in(0,1-c)$; if $\beta_{0}<\beta_{1}$, then $g<0$ for $a\in(0,1-c)$; and if $\beta_{0}=\beta_{1}$, then $g\equiv 0$ for $a\in(0,1-c)$.
Now we consider the case $\beta_{0}\beta_{1}>\lambda$. From (19), it is obvious that $g$ is monotone decreasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46). We now need to check whether $a^{*}\in[0,1-c]$. Notice that $a^{*}>1-c$ if and only if
$\beta_{0}-\beta_{1}>\frac{\beta_{0}\beta_{1}-\lambda}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c)),$ (47)
and $a^{*}<0$ if and only if
$\beta_{0}-\beta_{1}<-\frac{\beta_{0}\beta_{1}-\lambda}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c)),$ (48)
so we have the following cases: if (48) is satisfied, then $g<0$ for $a\in(0,1-c)$; if (47) is satisfied, then $g>0$ for $a\in(0,1-c)$; and if
$\beta_{0}-\beta_{1}\leq\left|\frac{\beta_{0}\beta_{1}-\lambda}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c))\right|,$ (49)
then $g>0$ for $a\in(0,a^{*})$ and $g<0$ for $a\in(a^{*},1-c)$.
Similarly we consider the case $\beta_{0}\beta_{1}<\lambda$. From (19), it is obvious that $g$ is monotone increasing on (0,1c)(0,1-c), and $g=0$ if and only if $a=a^{*}$ as defined in (46). We now need to check whether $a^{*}\in[0,1-c]$. Notice that $a^{*}>1-c$ if and only if
$\beta_{0}-\beta_{1}<-\frac{\lambda-\beta_{0}\beta_{1}}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c)),$ (50)
and $a^{*}<0$ if and only if
$\beta_{0}-\beta_{1}>\frac{\lambda-\beta_{0}\beta_{1}}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c)),$ (51)
so we have the following cases: if (50) is satisfied, then $g<0$ for $a\in(0,1-c)$; if (51) is satisfied, then $g>0$ for $a\in(0,1-c)$; and if
$\beta_{0}-\beta_{1}\leq\left|\frac{\lambda-\beta_{0}\beta_{1}}{\sqrt{\lambda}}\tanh(\sqrt{\lambda}(1-c))\right|,$ (52)
then $g<0$ for $a\in(0,a^{*})$ and $g>0$ for $a\in(a^{*},1-c)$.
This completes the proof together with Proposition 9. ∎
Remark on homogeneous Robin boundary
In [2], when studying the same problem under homogeneous Robin boundary ($\beta_{0}=\beta_{1}=\beta$), it is possible to reduce the condition $g=0$ to $\lambda=\beta_{0}\beta_{1}$, which could be plugged into $f=0$ to eliminate the terms involving $a$, which yields an equation of $\beta$, whose smallest positive solution can be expressed explicitly as $\beta^{*}=\beta^{*}(c,\kappa)$. Hence the final result in one dimension is given in cases where the criteria do not involve $\lambda$ which is a priori unknown. However, under our general inhomogeneous Robin boundary conditions, this is unfortunately impossible, since the condition $g=0$ does not help much in eliminating $a$ from the equation $f=0$ (cf. (19)) and (17). Nevertheless, in the next section we shall see that the criteria in Theorem 11 involving $\lambda$ could still be tested and thus the theorem itself be verified numerically.
3 Transfer–matrix shooting
In this section, we use a classical shooting method to numerically verify Theorem 11 for $\beta_{0}\beta_{1}\neq\lambda$.
For given $(c,\kappa,\beta_{0},\beta_{1})$ and a placement $a\in[0,1-c]$ of the favourable subinterval $(a,a+c)$, the equation
$\begin{cases}u^{\prime\prime}+\lambda\,m(x)\,u=0,\\ u^{\prime}(0)-\beta_{0}u(0)=0,\\ u^{\prime}(1)+\beta_{1}u(1)=0,\end{cases}$ (53)
with
$m(x)=\begin{cases}\kappa,&\text{ if }x\in(a,a+c),\\ -1,&\text{ otherwise,}\end{cases}$ (54)
admits explicit fundamental solutions on each subinterval:
$\begin{cases}u(x)=\rho\cosh(\sqrt{\lambda}\,x)+\sigma\sinh(\sqrt{\lambda}\,x),&x\in(0,a)\cup(a+c,1),\\ u(x)=\mu\cos(\sqrt{\lambda\kappa}\,x)+\nu\sin(\sqrt{\lambda\kappa}\,x),&x\in(a,a+c).\end{cases}$
Writing the state vector as $w=(u,u^{\prime})^{\top}$, then
$w^{\prime}(x)=(u(x)^{\prime},u(x)^{\prime\prime})^{\top}=\begin{pmatrix}0&1\\ -\lambda m(x)&0\end{pmatrix}w(x)=:Qw(x),$ (55)
where $m(x)\in\{-1,\kappa\}$ is piecewise constant. Now we integrate within an interval $x_{0},x]$ where $m(x)$ is constant, and get
$w(x)=e^{(x-x_{0})Q}w(x_{0}).$ (56)
The continuity of $u$ and $u^{\prime}$ at $x=a$ and $x=a+c$ thus yields
$w(1)=e^{(1-a-c)Q_{-1}}e^{cQ_{\kappa}}e^{aQ_{-1}}w(0),$ (57)
where for any $m\in\mathbb{R}$, $Q_{m}:=\begin{pmatrix}0&1\\ -\lambda m&0\end{pmatrix}$. Now that (53) is linear homogeneous with respect to $u$, we assume, without loss of generality that $u(0)=1$, then the Robin boundary condition gives $w(0)=(1,\beta_{0})^{\top}$ and $e_{2}w(1)+\beta_{1}e_{1}w(1)=0$, where $e_{1}=(1,0)$ and $e_{2}=(0,1)$. We now apply (57) and get
$R(a,\lambda):=e_{2}M(a,\lambda)\,(1,\beta_{0})^{\top}+\beta_{1}\,e_{1}M(a,\lambda)\,(1,\beta_{0})^{\top}=0,$ (58)
where $M(a,\lambda):=e^{(1-a-c)Q_{-1}}e^{cQ_{\kappa}}e^{aQ_{-1}}$. The principal eigenvalue we are interested in is interpreted as the smallest positive root of (58). We search for such roots on the window $\lambda\in\left(0,\frac{\pi^{2}}{4c^{2}\kappa}\right)$ using bisection on subintervals bracketed by a uniform scan in $\lambda$.
First, we pick a safe bound for $\lambda$ by setting
$\lambda_{\min}:=\max\left\{10^{-12},10^{-6}\frac{\pi^{2}}{4c^{2}\kappa}\right\},\ \ \ \lambda_{\max}:=(1-10^{-9})\frac{\pi^{2}}{4c^{2}\kappa},$ (59)
and take a uniform grid of sample $\lambda$-values:
$\lambda_{j}=\lambda_{\min}+j\frac{\lambda_{\max}-\lambda_{\min}}{N_{\lambda}},\ 0\leq j\leq N_{\lambda},$ (60)
where $N_{\lambda}$ is the number of subintervals we choose. Then a root of (58) must lie in a subinterval $(\lambda_{j},\lambda_{j+1})$ satisfying
$R(a,\lambda_{j})R(a,\lambda_{j+1})<0.$ (61)
If no such subinterval exists, we must refine the grid by increasing $N_{\lambda}$. Alternatively, we could lower $\lambda_{\min}$ to make it closer to 0.
Once a scan finishes with at least one subinterval where $R(a,\lambda)$ changes sign, we take the leftmost such subinterval $(\lambda_{j},\lambda_{j+1})$ and approach the eigenvalue $\lambda\in(\lambda_{j},\lambda_{j+1})$ by bisecting this subinterval. In other words, we check the sign of $\frac{\lambda_{j}+\lambda_{j+1}}{2}$ and refine the subinterval by its subinterval where $R(a,\lambda)$ changes sign, repeatedly until the length of the subinterval is smaller than a chosen tolerance, and output the eigenvalue as the midpoint of this final subinterval.
Finally, such numerical computation is done for a finite grid of $a$-values in the interval [0,1c][0,1-c], with a chosen reasonable resolution $N_{a}$, in order to produce the curve $a\mapsto\lambda(a)$ and to determine where the eigenvalue is minimised.
To compare the numerical result with Theorem 11 (non-degenerate cases $\beta_{0}\beta_{1}\neq\lambda$ only), we run the computation for a range of values for $(\beta_{0},\beta_{1})$ and for each pair of the Robin parameters, we verify the criteria in Theorem 11 for all computed eigenvalues $\lambda(a_{j})$, where $\{a_{j}\}_{j}$ are the grid of $a$-values in the interval [0,1c][0,1-c]. If a certain criterion in Theorem 11 is satisfied by all such $\lambda(a_{j})$, we classify the pair $(\beta_{0},\beta_{1})$ in the case corresponding to the valid criterion in Theorem 11, otherwise we do not classify this pair. For all those pairs of Robin parameters that are classified in a certain case of Theorem 11, we compare the predicted location for the minimum eigenvalue and the numerical result, and thus verify the correctness of Theorem 11.
Choice of parameters
We fix $N_{a}=81$, $N_{\lambda}=900$, and the tolerance $\texttt{tol}=10^{-10}$ of error. For Robin parameters, we sweep a rectangular grid $(\beta_{0},\beta_{1})\in[\beta_{\min},\beta_{\max}]^{2}$, with default $\beta_{\min}=0.2$, $\beta_{\max}=8.0$, and we take $n_{\beta}=10$ so that $n_{\beta}^{2}=100$ pairs are tested. The upper bound $\beta_{\max}$ is chosen to ensure that $\beta_{0}\beta_{1}-\lambda$ changes sign for the chosen $\kappa=2.0$ and $c=0.3$. In general, if some case from Theorem 11 is not found in a given $\beta$–grid, it is appropriate to widen $\beta_{\min},\beta_{\max}]$ or increase $n_{\beta}$.
Outputs
Our script produces:
  1. 1.
    a CSV file collecting all tested pairs and the classification, of which we report a compact excerpt in Table 1; and
  2. 2.
    up to six figures, one per case in Theorem 11, where we show in Figure 1.
In all cases, as Table 1 and Figure 1 clearly show, the prediction by Theorem 11 is accurate.
$\beta_{0}\beta_{1}>\lambda$
, $\beta_{1}\gg\beta_{0}$ (left).
$\beta_{0}\beta_{1}>\lambda$
, $|\beta_{0}-\beta_{1}|$ small (interior).
$\beta_{0}\beta_{1}>\lambda$
, $\beta_{0}\gg\beta_{1}$ (right).
$\beta_{0}\beta_{1}<\lambda$
, $\beta_{1}\gg\beta_{0}$ (left).
$\beta_{0}\beta_{1}<\lambda$
, $|\beta_{0}-\beta_{1}|$ small (either).
$\beta_{0}\beta_{1}<\lambda$
, $\beta_{0}\gg\beta_{1}$ (right).
Figure 1: Canonical examples for the six non-degenerate cases of Theorem 11.
Table 1: Excerpt (20 rows) from the batch verification on the grid of $(\beta_{0},\beta_{1})$.
$c$ $\kappa$ $\beta_{0}$ $\beta_{1}$ regime subcase predicted numeric comparison $\arg(\lambda_{\min})$
0.300 2.000 8.00 0.20 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 8.00 0.98 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 8.00 1.58 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 0.98 8.00 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 1.58 8.00 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 0.20 8.00 $\beta_{0}\beta_{1}>\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 5.64 2.00 $\beta_{0}\beta_{1}>\lambda$ $|\beta_{0}-\beta_{1}|$ small interior interior True 0.480
0.300 2.000 4.00 4.00 $\beta_{0}\beta_{1}>\lambda$ $|\beta_{0}-\beta_{1}|$ small interior interior True 0.480
0.300 2.000 2.00 5.64 $\beta_{0}\beta_{1}>\lambda$ $|\beta_{0}-\beta_{1}|$ small interior interior True 0.480
0.300 2.000 0.20 0.20 $\beta_{0}\beta_{1}<\lambda$ $|\beta_{0}-\beta_{1}|$ small either right True 0.000
0.300 2.000 0.20 0.98 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 0.20 1.58 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 0.20 2.38 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\ll\beta_{1}$ left left True 0.000
0.300 2.000 0.98 0.20 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 1.58 0.20 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 2.38 0.20 $\beta_{0}\beta_{1}<\lambda$ $\beta_{0}\gg\beta_{1}$ right right True 0.700
0.300 2.000 0.60 0.60 $\beta_{0}\beta_{1}<\lambda$ $|\beta_{0}-\beta_{1}|$ small either right True 0.000
0.300 2.000 0.98 0.98 $\beta_{0}\beta_{1}<\lambda$ $|\beta_{0}-\beta_{1}|$ small either right True 0.000
0.300 2.000 1.58 1.58 $\beta_{0}\beta_{1}<\lambda$ $|\beta_{0}-\beta_{1}|$ small either right True 0.000
0.300 2.000 2.38 2.38 $\beta_{0}\beta_{1}>\lambda$ $|\beta_{0}-\beta_{1}|$ small interior interior True 0.480
[…]
Acknowledgements
This work is partially supported by Grant No. SGS07/PŘF/2025, University of Ostrava.